OpenStudy (anonymous):

1/1+sin+1/1+cos+1/1sec+1/1+csc

OpenStudy (anonymous):

\[\frac{ 1 }{\sin x+1 }+\frac{ 1 }{ \cos x+1 }+\frac{ 1 }{\sec x+1 }+\frac{ 1 }{\csc x+1 }\] \[\frac{ 1 }{ \sec x+1 }=\frac{ 1 }{ \frac{ 1 }{\cos x }+1 }=\frac{ \cos x }{1+\cos x }\] \[\frac{ 1 }{ \csc x+1 }=\frac{ 1 }{ \frac{ 1 }{\sin x }+1 }=\frac{ \sin x }{ 1+\sin x }\] Now you can solve.